OpenStudy (anonymous):

Betty is having invitations printed for a holiday party. It costs $3.75 to set up the printing machine and $3.42 per invitation. How much will it cost Betty to print 150 invitations? $565.92 $153.75 $520.50 $516.75

OpenStudy (anonymous):

3.75 + 3.42(150) = 3.75 + 513 = $516.75

OpenStudy (anonymous):

thanks